OpenStudy (anonymous):

How do I solve (cos(pi/3 +h) - 1/2)/ h as h --> 0 but without lhopitals rule

OpenStudy (owlcoffee):

try using a trig. identity

OpenStudy (anonymous):

how would i do that ?

OpenStudy (anonymous):

first expand cos term.

OpenStudy (anonymous):

i have to use cos a plus b +from precalc) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

still stuck . i cant di anything with the minus 1/2 its in my way

OpenStudy (anonymous):

Are you trying to calculate a derivative?

OpenStudy (anonymous):

\[ \cos(x+y)=\cos(x)\cos(y)-\sin(x)\sin(y) \]